shirin's father is four times as old as she is. In 18 years he will be exactly twice as old as she. Finf ther present age.
Answer by kggoyani(2)

Let we assume that shirin is X years old, and her father is Y years old.
At, present the equation is y=4x......(1)
after 18 years age of shirin and her father will increase by 18 years, hence we can write an equation after 18 years is (y+18)=2(x+18)......(2)
now put y=4x in equation (y+18)=2(x+18)
so, (4x+18)=2(x+18)
4x+18=2x+36
4x-2x=36-18
2x=18
x=9
here, x=9 put in an equation (1).
so, y=4x
y=4(9)
y=36
we have found out that shirins's current age is 9 years and her father's current age is 36 years.